Grade 12th passAlgebra

Find ten rational numbers between 3/5 and ¾, ….......

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us make the denominators same, say 80.
3/5 = (3 × 16)/(5× 16) = 48/80
3/4=(3×20)/(4×20)=60/80
Ten rational numbers between 3/5 and 3⁄4 = ten rational numbers between 48/80 and 60/80 Therefore, ten rational numbers between 48/80 and 60/80 = 49/80, 50/80, 51/80, 52/80, 54/80, 55/80, 56/80, 57/80, 58/80, 59/80

Thanks